Sunspot Beverages, Ltd., of Fiji uses the weighted average method in its process costing system. It makes blended tropical frRequired 1 Required 2 Required 3 Required 4 Required 5 Calculate the Blending Departments equivalent units of production forRequired 1 Required 2 Required 3 Required 4 Required 5 Calculate the Blending Departments cost per equivalent unit for materCalculate the Blending Departments cost of ending work in process inventory for materials, conversion, and in total for JuneCalculate the Blending Departments cost of units transferred out to the Bottling Department for materials, conversion, and iPrepare a cost reconciliation report for the Blending Department for June. (Round your intermediate calculations to 2 decimal
Answer #1

1)

Equivalent Units of Production- Weighted Average Method
Units % Materials EUP- Materials % Conversion EUP- Conversion
Units completed and transferred out 160000 100% 160000 100% 160000
Units in ending work in process 40000 100% 40000 25% (40000*25%)= 10000
Total equivalent units of production 200000 200000 170000

2) Cost per equivalent unit= (Cost of Work in process, beginning+Cost added during June)/Equivalent units of production

Materials= $(25200+334800)/200000= $1.80

Conversion= $(24800+238700)/170000= $1.55

Materials Conversion
Cost per equivalent unit $1.80 $1.55

3) Cost of ending work in process of Materials= Units in ending work in process of materials*Cost per equivalent unit

= 40000*$1.80= $72000

Cost of ending work in process of Conversion= Units in ending work in process of Conversion*Cost per equivalent unit

= 10000*$1.55= $15500

Total Cost of ending work in process= $72000+15500= $87500

Materials Conversion Total
Cost of ending work in process $72000 $15500 $87500

4) Cost of units completed and transferred out

Materials= Units completed and transferred out*Cost per equivalent unit of materials

= 160000*$1.80= $288000

Conversion= Units completed and transferred out*Cost per equivalent unit of conversion

= 160000*$1.55= $248000

Total cost of units completed and transferred out= $288000+248000= $536000

Materials Conversion Total
Cost of units completed and transferred out $288000 $248000 $536000

5)

Blending Department
Cost Reconciliation
Cost to be accounted for:
Cost of beginning work in process (25200+24800) $50000
Cost added during the period (334800+238700) 573500
Total cost to be accounted for $623500
Cost accounted for as follows:
Cost of units completed and transferred out 536000
Cost of ending work in process 87500
Total cost accounted for $623500